OpenStudy (anonymous):

A. What is the slope of the line that passes through the points (4, 0) (0, 5)? B. Write a linear equation for the line.

OpenStudy (anonymous):

@Forte11 I suggest you try to answer it first using this formula and then I'll check afterwards. :) \[m=\frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]

OpenStudy (anonymous):

567945666667

OpenStudy (anonymous):

Is it -5 over 4?

OpenStudy (anonymous):

@Forte11 Yeah! Great job! :)

OpenStudy (anonymous):

How do I find out the linear equation for it? and thank you!

OpenStudy (anonymous):

That's easy. All you need to do is replace the missing variables in: y=mx+b You're most welcome. :)

OpenStudy (anonymous):

Thank you SO much!

OpenStudy (anonymous):

Anytime. :)

OpenStudy (anonymous):

Just checking its 5=m4+b?

OpenStudy (anonymous):

5 = \[5=(\frac{ -5 }{ 4 }) 0 + b \] And then, derive from there. :)

OpenStudy (anonymous):

What is B? Do I just leave it as b?

OpenStudy (anonymous):

No. You have to find it using the formula I filled for you. B will be equals to 5 in there. :)